The end of a flower box forms the shape of a trapezoid. The area of the trapezoid is 300 square centimeters. The bases are 16 centimeters and 24 centimeters in length. Find the height
Answer by nyc_function(2741)

We need the formula A = h/2(base 1 + base 2)
300 = h/2(16 + 24)
300 = h/2(40)
300 = 20h
300/20 = h
15 centimeters = h